OpenStudy (anonymous):

what is the answer to this question: two-fifths of the audience in a theatre are men, seven tenths of the remainder are women and the rest are children. There are 60 more women then children. How many people are there in the theatre altogehter

OpenStudy (anonymous):

M = 0.4 A W = 0.7(1-0.4)A = 0.42 A C = [1 - 0.4 -0.42] A = .18 A W = C + 60 or C = W - 60 A = M + W + C. W = 0.42 A = 0.18 A + 60 60 = 0.24 A A = 250 check M = 0.4 x 250 = 100 W = .42 A = 105 C = .18 A = 45 total is 250 and W = C+ 60, checks.

OpenStudy (anonymous):

thank you